# SWE-Synth: Synthesizing Verifiable Bug-Fix Data to Enable Large Language Models in Resolving Real-World Bugs

### v1

@@ Abstract
11footnotetext: *Co-author with equal contribution22footnotetext: Work in progress
@@ 1 Introduction
Large language models (LLMs) have made significant strides in automating software engineering (SE) tasks, from code generation Wei et al. (2023); Chen et al. (2021); Wang et al. (2023b); Li et al. (2022b); Zhuo et al. (2024); Bui et al. (2023); Manh et al. (2023); To et al. (2023) to bug fixing Jimenez et al. (2023); Xia et al. (2024a) and test synthesis Chen et al. (2022). Recent efforts have explored the development of LLM-based agents that can autonomously navigate repositories, localize bugs, generate patches, and verify correctness Yang et al. (2024); Jain et al. (2024); Liu et al. (2024); Phan et al. (2024); Nguyen et al. (2024). These agent-based approaches promise to scale software engineering workflows, but they require training data that is both high-quality and task-aligned.
Benchmarks like SWE-Bench Jimenez et al. (2023) and SWE-Bench Lite Yang et al. (2024) have emerged as standard evaluations for assessing autonomous software engineering (SWE) agents on realistic tasks such as GitHub issue resolution. These benchmarks have revealed a growing performance gap between proprietary and open-source models Xie et al. (2025); Jaech et al. (2024). While commercial agents backed by closed models (e.g., GPT-4, Claude 3.5) achieve strong performance Anthropic (2025), open models still lag behind, in part due to a lack of high-quality training environments and scalable datasets.
Unfortunately, existing datasets often fall short. Mined bug-fix pairs from open-source repositories Just et al. (2014); Herzig et al. (2013) are noisy, lack intermediate reasoning traces, and rarely include test-based verifiability Yan et al. (2024); Yang et al. (2024). Even curated benchmarks like SWE-Bench contain gold patches but lack comprehensive, executable environments for training. SWE-Gym Pan et al. (2024a), for example, partially addresses this by offering environments for a limited number of real issues, but its reliance on human-curated test cases limits its scalability.
Synthetic data generation offers a promising, scalable alternative Long et al. (2024); Wang et al. (2024a), yet existing techniques have their own shortcomings. Mutation-based methods Jia and Harman (2011) often yield trivial or unrealistic bugs. LLM-based synthetic data generation, while successful in instruction tuning Peng et al. (2023); Li et al. (2023), is not yet mature for complex SE tasks like bug repair, fault localization, or test-driven development. Furthermore, synthetic data frequently lacks contextual grounding or verification, limiting its usefulness for downstream models Ding et al. (2023); Pan et al. (2024a).
@@
  • Automated verification: Each buggy example should be paired with test cases that verify both the existence of the bug and the correctness of the fix Le et al. (2022); Li et al. (2025); Zeng et al. (2025); Chu et al. (2025).
  • Repair traces: Intermediate steps—such as localization, fixing, and refinement—should be included to model the developer’s iterative reasoning process Yang et al. (2024); stechly2024chain; Zhou et al. (2024b).
In this work, we propose SWE-Synth, a scalable framework for generating high-fidelity, verifiable synthetic bug-fix datasets using LLM agents that simulate human debugging workflows. Our pipeline produces not only buggy-fix pairs, but also structured intermediate traces and test logs that enable modern learning algorithms such as reinforcement learning and rejection sampling Liu et al. (2024); Wei et al. (2025); Yao et al. (2022). In summary, we make the following contributions:
@@ 2 Illustrating Example
This section illustrates how LLMs could be used to generate buggy code. In Figure 1, we present an example on the pydantic library. The original code defines the Base64Encoder class, which contains three class methods: decode, encode, and get_json_format Pydantic . We begin by masking the implementation details of these methods and then use LLM to regenerate them (Section 4 for details). The resulting implementation often deviates from the original code, producing different variants, e.g., $v_{1}$ and $v_{2}$ (Figure 1).
@@
In the second variant $v_{2}$, the LLM’s implementation replaces encodebytes and decodebytes with b64encode and b64decode from the base64 library, deviating from the original code. This change causes the test test_construct_nested to fail due to subtle output differences between b64encode and encodebytes Fati . This error exemplifies a typical API misuse, where an incorrect API is selected Amann et al. (2019), often due to a lack of domain knowledge about the base64 library. Moreover, this LLM-generated error demonstrates the potential for introducing non-trivial bugs that require fixes across multiple locations—reflecting the complexity of real-world debugs. Traditional mutation-based bug synthesis struggles to generate such multi-location bugs due to the combinatorial explosion of mutations, making LLM-based solutions a promising alternative.
@@ 4 LLM-based Variant Synthesis
INPUT: Original and correct program $p_{o}$, set of components $\mathcal{C}$ in program $p_{o}$, large language model $LLM$
@@ 4.1 Component Selection
In the component selection phase, our objective is to identify a component $c$ from the set of all components $\mathcal{C}$ within the original program $p_{o}$ as the target to re-implement. A component $c$ is broadly defined as any program entity within the program (codebase), such as variables, statements, functions, classes, files, or modules. The selection of $c$ is a pivotal decision in our procedure, as it determines the locus of the synthetic bug introduced into the program. This, in turn, influences the characteristics and significance of the resulting program variants. A poorly chosen component—such as a peripheral or infrequently utilized part of the program—often leads to the creation of trivial or irrelevant bugs that fail to challenge the program in meaningful and practical ways. Conversely, a strategically selected component, particularly one that is central to the program’s operation and well-covered by test cases, can generate variants that closely resemble realistic and impactful defects. We presented two strategies to select the component for rewriting:
Naive Strategy with Uniform Random Selection: A straightforward method for selecting $c$ is to sample it uniformly at random from $\mathcal{C}$: $c\sim\text{Uniform}(C)$. In this strategy, each component in $\mathcal{C}$ has an equal probability of selection, calculated as $P(c)=\frac{1}{|C|}$, where $|\mathcal{C}|$ denotes the total number of components.
While computationally simple, this strategy assumes all components are equally important to the program’s behavior—an assumption that typically does not hold. Real-world programs are heterogeneous, with some components (e.g., core functions) playing a central role, and others (e.g., rarely executed statements) being much less critical. Consequently, reimplementing a peripheral component can yield unchallenging and less valuable variants.
Test Coverage-Based Selection Strategy: To overcome the limitations of uniform random selection, we propose a test coverage-based selection strategy. This strategy prioritizes components based on the extent to which they are exercised by the test suite, reflecting their importance to the system’s functionality. We hypothesized that components that are covered by a larger number of test cases are thus likely to be central to the program’s core behaviors, making them prime candidates for rewriting. Moreover, the files/components with high test coverages would need more attention likely due to the history of buggy code in the past. By targeting these components, we can increase the likelihood of generating synthetic bugs that are impactful, and representative of real-world defects.
A practical challenge in implementing this strategy is the need to execute the test suite to obtain coverage data. Running all test cases can be time-consuming. To address this, we adopt a design that amortizes the cost of test execution across multiple potential mutations. Specifically, once we capture coverage data for a given repository commit, we reuse that same data to inform the selection of multiple mutation locations. This allows us to avoid repeated test coverage runs for each mutation, thereby reducing overhead.
@@ 4.2 LLM-based Re-implementations as Variants
{function_signature} … your code goes here …
@@ 6 SWE-Synth: A Synthetic Dataset Reflecting Real-World Bug-Fix Patterns
To enable supervised fine-tuning (SFT), we extracted ground-truth patches from successful rollout trajectories (Section 5). While our framework is compatible with various agent scaffolds, we used the open-source MoatlessTool Örwall (2024), which integrates code search with LLM-based retrieval. Each variant’s failed test log was truncated to 16k tokens and included in the input prompt. For embedding-based retrieval, we used Jina Embeddings v2 Base Code günther2024jinaembeddings28192token.
@@ 7.2.1 Models, Benchmarks, and Procedure
We chose the Qwen 2.5 Coder Instruct models Hui et al. (2024) with different sizes 7B, 14B, and 32B. After training the models with SWE-Synth and the manual dataset in Pan et al. (2024b), we evaluated them primarily on the SWE-Bench Lite Jimenez et al. (2024) dataset, which includes 300 bugs from 12 popular Python repositories, each having a patch based on its bug report. Additionally, we assess performance on the BugsInPyWidyasari et al. (2020) dataset, which contains 493 bugs from 17 projects, though we successfully reproduced only 317 bugs (due to missing required libraries). To align with our problem formulation for APR (see Section 3), we use failed test logs from these datasets as task prompts.
All training experiments were conducted using LoRA Hu et al. (2022) with a rank of 8, an alpha of 16, and a learning rate of $2\times 10^{-5}$. Training was performed on 4 H100 GPUs for 5 epochs with a global batch size of 8. We utilized the LLAMA-Factory framework Zheng et al. (2024) and incorporated LongLoRA Chen et al. (2023) to support a context length of 65,536 during training. Unless otherwise specified, we fine-tuned from Qwen 2.5 Coder Instruct Hui et al. (2024) with a model size of 14B. For the 32B model, we employed QLoRA Dettmers et al. (2023) for fine-tuning.
@@ 7.3 Manual Data versus Synthetic Data
[…]
@@ 7.3.1 Model performance comparison on synthetic and manual data
Table 3: APR Performance with 1K success trajectories (RQ1)
[…]
@@
SWE-Bench
Lite
11.33% 11.67% 2.33% 1.33% 33.00% 25.67%
@@
Same number of variants: For this setting, we fixed the sample size of 230 instances from SWE-Gym Lite and an equal number of variants from SWE-Synth. However, each dataset underwent three rounds of rollouts, with supervised fine-tuning performed on the success trajectories collected under different capping numbers of per-bug instances. As shown in Table 2, under the most lenient capping (Cap 3), the advantages of SWE-Synth become particularly clear. Although the overall resolve rate remains similar—12.00% for SWE-Gym versus 11.67% for SWE-Synth on SWE-Bench Lite—SWE-Synth demonstrates consistent superior performance on the BugsInPy benchmark on all metrics and caps (with most significant gap in Cap 1, 19.24% resolve rate compared to only 16.09%). This result suggests that SWE-Synth ’s training signal is richer, likely due to its controlled diversity and balance of both easy and challenging bugs via capping the numbers of per-bug instances.
Same total number of trajectories: In the second setting, we standardized the total training data by randomly sampling 1,000 success trajectories from each dataset, without enforcing constraints on the number of variants or per-bug instance capping. As seen in Table 3, the model performance on the two datasets converges, with that on synthetic data often matching—or even slightly surpassing the manually curated dataset in terms of resolve rate (SWE-Synth’s 11.67% vs 11.33% of manual data). For BugsInPy, the performance remains nearly identical across manual and synthetic datasets, with both achieving the same correctness rate (1.89%) and empty patch rate (28.08%) but slightly lags behind the manual dataset in terms of resolve rate (17.67% vs 19.56%). Essentially, while the manual dataset benefits from high-quality, real-world issues, the scalability (Section 7.3.2) and diversity of SWE-Synth allow it to compensate for its quality, leading to comparable performance.
@@ 7.4 Synthetic Data and Model Size
We aim to determine if smaller models benefit disproportionately from additional bug-fix data, or if larger models continue to see comparable gains despite their stronger inherent capabilities. We randomly sampled 1,027 success trajectories to fine-tune three Qwen 2.5 Coder models with 7B, 14B, and 32B parameters.
Table 5 reveals that the 7B model benefits most from fine-tuning, as demonstrated by a marked reduction in its empty patch rate and a significant increase in its resolve rate compared to its modest zero-shot performance. In larger models, the improvement gap between zero-shot and fine-tuned performance narrows, partly because the training trajectories generated by Qwen 2.5 Coder Instruct 32B (see Section 6) provide a strong teaching signal that disproportionately benefits smaller, student models via knowledge distillation with rejection sampling—a process described as Weak-to-Strong improvement in Bansal et al. (2024). Our results align with prior work Pan et al. (2024b); Bansal et al. (2024), which shows that the advantage of Weak-to-Strong improvement diminishes as the student model size increases. Overcoming this limitation may require a fundamental algorithmic change to enable self-improvement, where the teacher and student models are of equal size. One potential avenue is to employ Reinforcement Learning with Verifiable Rewards Guo et al. (2025), which allows the model to learn from its own mistakes—a strategy that is also feasible with SWE-Synth given its automated verifiable bugs and intermediate steps.
@@
[…]
@@ 7.5 In-depth Study of Data Synthesis Pipeline
[…]
@@ 7.5.2 Impact of component selection
We employ two component-selection strategies for rewriting: uniform random selection and test coverage-based selection (Section 4.1). To evaluate their impact, we fine-tuned on 400 success-rollout trajectories generated by each strategy. As seen in Table 7, coverage-based selection achieves a higher resolve rate (12.33% vs. 12.00%) and a lower empty patch rate (26.67% vs. 35.67%) compared to uniform sampling. This aligns with prior observations that trivial or irreproducible bugs often lead to more empty patches Yang et al. (2024). Moreover, coverage-based selection produces variants harder to fix, with 87.39% of them remaining unresolved by Qwen 2.5 Coder Instruct 32B + Moatless Tools (versus 84.09% for uniform selection). By focusing on components covered by more tests, the coverage-based strategy synthesizes a variety of realistic, non-trivial bugs.
@@
[…]

### v2

@@ Abstract
[…]
@@ 1 Introduction
Large language models (LLMs) have made significant strides in automating software engineering (SE) tasks, from code generation Wei et al. (2023); Chen et al. (2021); Wang et al. (2023); Li et al. (2022b); Zhuo et al. (2024); Bui et al. (2023); Manh et al. (2023); To et al. (2023) to bug fixing Jimenez et al. (2023); Xia et al. (2024a) and test synthesis Chen et al. (2022); He et al. (2025); Li et al. (2024). Recent efforts have explored the development of LLM-based agents that can autonomously navigate repositories, localize bugs, generate patches, and verify correctness Yang et al. (2024b); Jain et al. (2024); Liu et al. (2024a); Phan et al. (2024); Nguyen et al. (2024a). These agent-based approaches promise to scale software engineering workflows, but they require training data that is both high-quality and task-aligned.
Benchmarks like SWE-Bench Jimenez et al. (2023) and SWE-Bench Lite Yang et al. (2024b) have emerged as standard evaluations for assessing autonomous software engineering (SWE) agents on realistic tasks such as GitHub issue resolution. These benchmarks have revealed a growing performance gap between proprietary and open-source models Xie et al. (2025); Jaech et al. (2024). While commercial agents backed by closed models (e.g., GPT-4, Claude 3.5) achieve strong performance Anthropic (2025), open models still lag behind, in part due to a lack of high-quality training environments and scalable datasets. Unfortunately, existing datasets often fall short. Mined bug-fix pairs from open-source repositories Just et al. (2014); Herzig et al. (2013) are noisy, lack intermediate reasoning traces, and rarely include test-based verifiability Yan et al. (2024); Yang et al. (2024b). Even curated benchmarks like SWE-Bench contain gold patches but lack comprehensive, executable environments for training. SWE-Gym Pan et al. (2024a), for example, partially addresses this by offering environments for a limited number of real issues, but its reliance on human-curated test cases limits its scalability. Synthetic data generation offers a promising, scalable alternative Long et al. (2024); Wang et al. (2024b), yet existing techniques have their own shortcomings. Mutation-based methods Jia and Harman (2011) often yield trivial or unrealistic bugs. LLM-based synthetic data generation, while successful in instruction tuning Peng et al. (2023); Li et al. (2023), is not yet mature for complex SE tasks like bug repair, fault localization, or test-driven development. Furthermore, synthetic data frequently lacks contextual grounding or verification, limiting its usefulness for downstream models Ding et al. (2023); Pan et al. (2024a).
@@
  • Automated verification: Each buggy example should be paired with test cases that verify both the existence of the bug and the correctness of the fix Le et al. (2022); Li et al. (2025); Anonymous (2024); Chu et al. (2025).
  • Repair traces: Intermediate steps—such as localization, fixing, and refinement—should be included to model the developer’s iterative reasoning process Yang et al. (2024b); Stechly et al. (2024); Zhou et al. (2024).
In this work, we propose SWE-Synth, a scalable framework for generating high-fidelity, verifiable synthetic bug-fix datasets using LLM agents that simulate human debugging workflows. Our pipeline produces not only buggy-fix pairs, but also structured intermediate traces and test logs that enable modern learning algorithms such as reinforcement learning and rejection sampling Liu et al. (2024a); Wei et al. (2025); Yao et al. (2022). In summary, we make the following contributions:
@@ 2 Illustrating Example
This section illustrates how LLMs could be used to generate buggy code. In Figure 1, we present an example on the pydantic library. The original code defines the Base64Encoder class, which contains three class methods: decode, encode, and get_json_format Pydantic (2025). We begin by masking the implementation details of these methods and then use LLM to regenerate them (Section 4 for details). The resulting implementation often deviates from the original code, producing different variants, e.g., $v_{1}$ and $v_{2}$ (Figure 1).
@@
In the second variant $v_{2}$, the LLM’s implementation replaces encodebytes and decodebytes with b64encode and b64decode from the base64 library, deviating from the original code. This change causes the test test_construct_nested to fail due to subtle output differences between b64encode and encodebytes Fati (2025). This error exemplifies a typical API misuse, where an incorrect API is selected Amann et al. (2019), often due to a lack of domain knowledge about the base64 library. Moreover, this LLM-generated error demonstrates the potential for introducing non-trivial bugs that require fixes across multiple locations—reflecting the complexity of real-world debugs. Traditional mutation-based bug synthesis struggles to generate such multi-location bugs due to the combinatorial explosion of mutations, making LLM-based solutions a promising alternative.
@@ 4 LLM-based Variant Synthesis
INPUT: Original and correct program $p_{o}$, set of components $\mathcal{C}$ in program $p_{o}$, large language model LLMLLM
@@ 4.1 Component Selection
In the component selection phase, our objective is to identify a component $c\in\mathcal{C}$ from the original program $p_{o}$ as the target for re-implementation. A component $c$ may be any program entity—such as a variable, statement, function, class, file, or module. This choice is pivotal: it determines where the synthetic bug will be injected and strongly influences the quality of the resulting variants. Selecting peripheral or rarely executed components often leads to trivial bugs that do not meaningfully test the program, whereas targeting central components, especially those covered by tests, produces variants that resemble realistic and impactful defects. We examine two strategies:
Naive Uniform Selection: The simplest approach samples $c$ uniformly at random from $\mathcal{C}$, assigning $P(c)=\frac{1}{|\mathcal{C}|}$. This method is computationally cheap but assumes all components contribute equally to program behavior. In practice, programs are heterogeneous: some components (e.g., core utility functions) are critical, while others (e.g., corner-case statements) are rarely relevant. As a result, this strategy often yields less challenging bugs of limited practical value.
Test Coverage-Based Selection: To address this limitation, we prioritize components according to their test coverage. Components executed by more test cases are more central to program behavior and more likely to contain bugs historically, making them better candidates for rewriting. This strategy increases the chance of generating synthetic bugs that are both impactful and representative of real-world defects. A practical challenge is the overhead of collecting coverage data, which requires running the test suite. To mitigate this, we cache coverage information for each repository commit and reuse it across multiple mutations, amortizing the cost and enabling efficient large-scale bug generation.
@@ 4.2 LLM-based Re-implementations as Variants
{function_signature}
  … your code goes here …
@@ 6 SWE-Synth: A Synthetic Dataset Reflecting Real-World Bug-Fix Patterns
To enable supervised fine-tuning (SFT), we extracted ground-truth patches from successful rollout trajectories (Section 5). While our framework is compatible with various agent scaffolds, we used the open-source MoatlessTool Örwall (2024), which integrates code search with LLM-based retrieval. Each variant’s failed test log was truncated to 16k tokens and included in the input prompt. For embedding-based retrieval, we used Jina Embeddings v2 Base Code Günther et al. (2024).
@@ 7.2.1 Models, Benchmarks, and Procedure
We chose the Qwen 2.5 Coder Instruct models Hui et al. (2024) with different sizes 7B, 14B, and 32B. After training the models with SWE-Synth and the manual dataset in Pan et al. (2024b), we evaluated them primarily on the SWE-Bench Lite Jimenez et al. (2024) dataset, which includes 300 bugs from 12 popular Python repositories, each having a patch based on its bug report. We also evaluate on the SWE-Bench Verified Chowdhury et al. (2024) dataset, which includes 500 bugs from the same set of repositories, but with human validation to ensure that each issue contains sufficient information to be solvable. Additionally, we assess performance on the BugsInPy Widyasari et al. (2020) dataset, which contains 493 bugs from 17 projects, though we successfully reproduced only 317 bugs (due to missing required libraries). To align with our problem formulation for APR (see Section 3), we use failed test logs from these datasets as task prompts.
All training experiments were conducted using LoRA Hu et al. (2022) with a rank of 8, an alpha of 16, and a learning rate of $2\times 10^{-5}$. Training was performed on 4 H100 GPUs for 5 epochs with a global batch size of 8. We utilized the LLAMA-Factory framework Zheng et al. (2024) and incorporated LongLoRA Chen et al. (2023) to support a context length of 65,536 during training. Unless otherwise specified, we fine-tuned from Qwen 2.5 Coder Instruct Hui et al. (2024) with a model size of 14B.
@@ 7.3 Manual Data versus Synthetic Data
Dataset $\%$ Resolved $(\uparrow)$ $\%$ Correct $(\uparrow)$ $\%$ EP $(\downarrow)$
@@ 7.3.1 Model performance comparison on synthetic and manual data
Table 3: APR Performance of Qwen 2.5 Coder Instruct 14B fine-tuned with 1K success trajectories from 32B (RQ1)
Dataset $\%$ Resolved $(\uparrow)$ $\%$ Correct $(\uparrow)$ $\%$ EP $(\downarrow)$
@@
SWE-Bench Lite 11.33% 11.67% 2.33% 1.33% 33.00% 25.67%
SWE-Bench Verified 11.40% 11.00% 1.60% 1.80% 31.40% 28.80%
@@
Table 4: APR Performance of Qwen 2.5 Coder Instruct 32B fine-tuned with 2136 success trajectories from DeepSeek-v3-0324 (RQ1)
Dataset $\%$ Resolved $(\uparrow)$ $\%$ Correct $(\uparrow)$ $\%$ EP $(\downarrow)$
manual synth manual synth manual synth
SWE-Bench Lite 23.67% 24.33% 3.33% 2.00% 13.00% 5.00%
SWE-Bench Verified 18.20% 22.20% 3.80 % 3.20% 17.80% 8.00%
Same number of variants: We fixed the sample size to 230 instances from SWE-Gym Lite and an equal number from SWE-Synth, each undergoing three rollout rounds. Supervised fine-tuning was performed on success trajectories with varying caps on per-bug instances. As shown in Table 2, under the most lenient cap (Cap 3), SWE-Synth ’s advantage is clear. While overall resolve rates on SWE-Bench Lite are similar (12.00% vs. 11.67%), SWE-Synth consistently outperforms on BugsInPy, with the largest gap at Cap 1 (19.24% vs. 16.09%). This suggests SWE-Synth provides a richer training signal through controlled diversity and balanced bug difficulty.
Same total number of trajectories: Here, we standardized training data to 1,000 success trajectories per dataset, without capping or variant constraints. Table 3 shows the two datasets converge: SWE-Synth matches or slightly surpasses manual data on SWE-Bench Lite (11.67% vs. 11.33%). On BugsInPy, both achieve the same correctness (1.89%) and empty patch rates (28.08%), with manual data slightly ahead in resolve rate (19.56% vs. 17.67%). Thus, while manual data benefits from real-world quality, the scalability and diversity of SWE-Synth compensate for this, yielding comparable performance.
To further test rollout effects under fixed trajectory budgets, we used DeepSeek-v3-0324 DeepSeek-AI (2024) with $\texttt{CodeXEmbed}_{\text{2B}}$ Liu et al. (2024b) to sample 2,136 success trajectories per dataset, fine-tuning Qwen 2.5 Coder Instruct 32B (context length 32,768). The baseline resolve rate on SWE-Bench Verified was 16.4%. As shown in Table 4, training on SWE-Synth boosted this to 22.20%, clearly surpassing manual data (18.20%). This improvement stems from SWE-Synth ’s larger set of bug instances and trajectory diversity, suggesting the synthetic advantage grows with larger fine-tuning data and stronger teacher models.
@@ 7.4 Synthetic Data and Model Size
We investigate whether smaller models gain disproportionately from additional bug-fix data or if larger models still see comparable improvements. We fine-tuned Qwen 2.5 Coder models with 7B, 14B, and 32B parameters on 1,027 success trajectories, using QLoRA Dettmers et al. (2023) for the 32B model.
Table 6 shows that the 7B model benefits most, with a sharp drop in empty patch rate and a clear rise in resolve rate over its weak zero-shot baseline. For larger models, gains narrow since the training trajectories—generated by Qwen 2.5 Coder Instruct 32B—already provide a strong signal, effectively distilling knowledge from strong teacher to smaller students. This Weak-to-Strong effect Bansal et al. (2024) diminishes with increasing student size, consistent with prior work Pan et al. (2024b); Bansal et al. (2024). Addressing this may require self-improvement methods, such as Reinforcement Learning with Verifiable Rewards Guo et al. (2025), which SWE-Synth can support through its automated verifiable bugs and intermediate steps.
@@
Dataset $\%$ Resolved $(\uparrow)$ $\%$ Correct $(\uparrow)$ $\%$ EP $(\downarrow)$
@@ 7.5 In-depth Study of Data Synthesis Pipeline
Level of granularity Difficulty $(\%,\uparrow)$ Trained model performance
$\%$ R $(\uparrow)$ $\%$ C $(\uparrow)$ $\%$ EP $(\downarrow)$
@@ 7.5.2 Impact of component selection
We employ two component-selection strategies for rewriting: uniform random selection and test coverage-based selection (Section 4.1). To evaluate their impact, we fine-tuned on 400 success-rollout trajectories generated by each strategy. As seen in Table 8, coverage-based selection achieves a higher resolve rate (12.33% vs. 12.00%) and a lower empty patch rate (26.67% vs. 35.67%) compared to uniform sampling. This aligns with prior observations that trivial or irreproducible bugs often lead to more empty patches Yang et al. (2024b). Moreover, coverage-based selection produces variants harder to fix, with 87.39% of them remaining unresolved by Qwen 2.5 Coder Instruct 32B + Moatless Tools (versus 84.09% for uniform selection). By focusing on components covered by more tests, the coverage-based strategy synthesizes a variety of realistic, non-trivial bugs.
@@
Component selection strategy Difficulty $(\%,\uparrow)$ Trained model performance
$\%$ R $(\uparrow)$ $\%$ C $(\uparrow)$ $\%$ EP $(\downarrow)$